Arturo is three times as old as his son. In five years, Arturo will be two and a half times as old as his son. What are their ages?

Answer by ankor@dixie-net.com(22740)

Arturo is three times as old as his son.
a = 3s
:
In five years, Arturo will be two and a half times as old as his son.
a + 5 = 2.5(s+5)
a + 5 = 2.5s + 12.5
a = 2.5s + 12.5 - 5
a = 2.5s + 7.5
Replaces a with 3s and solve


What are their ages?